# Physics-informed Reduced Order Modeling of Time-dependent PDEs via Differentiable Solvers

### v1

@@ Abstract
Reduced-order modeling (ROM) of time-dependent and parameterized differential equations aims to accelerate the simulation of complex high-dimensional systems by learning a compact latent manifold representation that captures the characteristics of the solution fields and their time-dependent dynamics. Although high-fidelity numerical solvers generate the training datasets, they have thus far been excluded from the training process, causing the learned latent dynamics to drift away from the discretized governing physics. This mismatch often limits generalization and forecasting capabilities. In this work, we propose Physics-informed ROM ($\Phi$-ROM) by incorporating differentiable PDE solvers into the training procedure. Specifically, the latent space dynamics and its dependence on PDE parameters are shaped directly by the governing physics encoded in the solver, ensuring a strong correspondence between the full and reduced systems. Our model outperforms state-of-the-art data-driven ROMs and other physics-informed strategies by accurately generalizing to new dynamics arising from unseen parameters, enabling long-term forecasting beyond the training horizon, maintaining continuity in both time and space, and reducing the data cost. Furthermore, $\Phi$-ROM learns to recover and forecast the solution fields even when trained or evaluated with sparse and irregular observations of the fields, providing a flexible framework for field reconstruction and data assimilation. We demonstrate the framework’s robustness across different PDE solvers and highlight its broad applicability by providing an open-source JAX implementation readily extensible to other PDE systems and differentiable solvers.
@@ 2.3 Reduced-ordered temporal evolution
[…]
@@ 3.1 Training objective
In Eq. (10), $\lambda$ controls the regularization effect of the dynamics loss. Choice of $\lambda$ depends on the sensitivity of $\mathcal{S}$ to the noise and errors in the reconstructed fields. We chose $\lambda=0.5-0.8$ depending on the solvers (see Appendix F.3). See Appendix B for the detailed training procedure.
@@ 4.1 Experimental settings
For each problem, models are trained on $M_{tr}$ trajectories of $T_{tr}$ time steps and evaluated on $M_{te}$ trajectories to assess forecasting and generalization performance. Evaluation is performed by forecasting solutions for unseen initial conditions or parameters in $\mathbf{U}_{te}$ for $T_{tr}$ (interpolation) and $T_{te}$ (extrapolation) time steps. Among the problems, Burgers’ and LBM are parameterized by their source term and Reynolds number, respectively, while the others vary in their initial conditions. Please refer to Appendix F for a detailed description of our experimental settings and Appendix G for extended results.
@@ 4.3 Forecasting performance on full grid
Full training $\mathcal{X}_{tr}=\mathcal{X}_{\mathcal{S}}=\mathcal{X}_{te}$
@@
Sparse training $|\mathcal{X}_{tr}|=5\%|\mathcal{X}_{\mathcal{S}}|,\,\mathcal{X}_{te}=\mathcal{X}_{\mathcal{S}}$
@@ 5 Conclusion
We introduced $\Phi$-ROM, a physics-informed ROM that enforces the governing physics directly in the reduced latent space by learning the dynamics directly from a differentiable PDE solver. We showed that compared to other physics-informed strategies and data-driven methods for constructing ROMs, $\Phi$-ROM generalizes better to unseen dynamics induced by new initial conditions or parameters, forecasts beyond the training time horizon, and recovers the full solution fields when trained or tested with partial observations. Furthermore, $\Phi$-ROM evolves the solution in the reduced latent space for accelerated forecasting while being mesh-free and continuous in space and time.
We acknowledge that requiring access to differentiable PDE solvers limits $\Phi$-ROM’s immediate applicability, although ongoing developments in differentiable programming offer promising avenues. The inclusion of the PDE solver also increases the training cost compared to purely data-driven methods. However, considering the superior data efficiency of $\Phi$-ROM, we believe this one-time training cost is justified. Additionally, although $\Phi$-ROM was successfully trained with sparse data, extending the framework to data assimilation settings with limited sensor measurements is a compelling area for further exploration. Finally, this study focused on PDEs with first-order time derivatives, and extending to higher-order systems (e.g. wave equations) is left for future research.
@@ Appendix A Related works
Our proposed framework of $\Phi$-ROM adopts a similar setup to DINo [13], but benefits from a physics-informed training strategy. $\Phi$-ROM directly trains the dynamics network with time derivatives of the reconstructed fields, obtained from a differentiable solver, in contrast to DINo, which relies on time-integration during training using neural ODEs to evolve the latent coordinates. We also leverage hyperreduction [22] for fast and efficient training of the dynamics network in $\Phi$-ROM, unlike other similar approaches that also rely on the time derivatives of the reconstructed fields [12, 19]. Moreover, $\Phi$-ROM differs substantially from CROM in the sense that it enforces the governing physics in the latent space and evolves the solution directly in the latent space as opposed to the full physical space.
@@ B.1 Training procedure
2:$\Gamma\leftarrow\left\{\alpha_{i}^{t}=\mathbf{0}\mid i\in\{0,\dots,M_{tr}\},\,t\in\{0,\dots,T_{tr}\}\right\}$
@@ B.2 Dynamics loss and hyper-reduction
2:     $\hat{\mathbf{u}}\leftarrow D_{\theta}(\alpha,\mathcal{X}_{\mathcal{S}})$triangleright Reconstruct the field on the solver grid ($\hat{\mathbf{u}}\in\mathbb{R}^{N\times m},N=\left|X_{\mathcal{S}}\right|$)
3:     $\dot{\hat{\mathbf{u}}}\leftarrow(\hat{\mathbf{u}},\mathcal{S})$triangleright Compute time derivative, see Appendix B.4
4:     $\mathcal{X}^{\downarrow}_{\mathcal{S}}\sim\mathcal{X}_{\mathcal{S}}$triangleright Randomly sample $\gamma N$ coordinates (where $N=|\mathcal{X}_{\mathcal{S}}|$)
5:     $\dot{\hat{\mathbf{u}}}^{\downarrow}\leftarrow\dot{\hat{\mathbf{u}}}|_{\mathcal{X}^{\downarrow}_{\mathcal{S}}}$triangleright Sample time derivatives at $\mathcal{X}_{\mathcal{S}}^{\downarrow}$ and flatten ($\dot{\hat{\mathbf{u}}}^{\downarrow}\in\mathbb{R}^{\gamma Nm}$)
6:     $J^{\downarrow}\leftarrow\nabla_{\alpha}D_{\theta}(\alpha,\mathcal{X}^{\downarrow}_{\mathcal{S}})$triangleright Jacobian via forward mode AD ($J^{\downarrow}\in\mathbb{R}^{\gamma Nm\times k}$)
7:     $\dot{\alpha}^{*}\leftarrow\operatorname*{\arg\!\min}_{\dot{\alpha}}\left\|\dot{\hat{\mathbf{u}}}^{\downarrow}-J^{\downarrow}\dot{\alpha}\right\|$triangleright Solve least-squares ($\dot{\alpha}^{*}\in\mathbb{R}^{k}$)
8:     return $\ell(\Psi_{\phi}(\alpha),\dot{\alpha}^{*})$triangleright Return dynamics loss
@@ Appendix C Inference details
2:$\hat{\alpha}^{t_{0}}\leftarrow\mathbf{0}$
3:for each inversion step $\triangleright$ Inversion ($D_{\theta}^{\dagger}(\mathbf{u}^{t_{0}},\mathcal{X})$)
@@
7:$\tilde{\alpha}^{T_{*}}\leftarrow ODESolve(\hat{\alpha}^{t_{0}},t_{0},T_{*})$triangleright Integrating the latent coordinates
8:$\tilde{\mathbf{u}}^{T_{*}}\leftarrow D_{\theta}(\tilde{\alpha}^{T_{*}},\mathcal{X}^{\prime})$triangleright Decoding the final state
@@ Appendix D Detailed description of the datasets
Diffusion $100$ $32$ $16$ $25$ $200$ $42\times 42$ $1$
Burgers’ $8$ $9$ $100$ $200$ $256$ $1$
N-S $256$ $64$ $16$ $50$ $200$ $64\times 64$ $2$
KdV $512$ $64$ $16$ $40$ $80$ $64\times 64$ $1$
LBM $40$ $40+40$ $4$ $100$ $125$ $180\times 82$ $2$
@@ D.5 Lattice Boltzmann equation
We used XLB in 2D to simulate flow around a circular cylinder with the D2Q9 lattice, BGK collision operator, and suitable boundary conditions representing the flow inlet and outlet, and the no-slip boundaries for a set of Reynolds numbers, $Re=UD/\nu$ where $U$ is the mean speed of the incoming flow, $D$ is the diameter of the cylinder, and $\nu$ is the kinematic viscosity. For sufficiently large Reynolds numbers, this configuration leads to the formation of the Kármán vortex street caused by vortex shedding. We generate the data on a $180\times 82$ grid with Reynolds numbers ranging from $100$ to $200$ and simulate for $4000$ solver steps. We discard the data from the first $3000$ steps of the solver when the flow is still forming. Furthermore, since LBM makes very small steps, we save the remaining simulated data only every $5$ steps. As such, each trajectory has $200$ time steps, where $T_{tr}=100$ are used for training and the rest are used for testing. However, we report the time extrapolation errors in the paper for only $T_{te}=125$ as both $\Phi$-ROM and DINo fail to capture the long-horizon dynamics of LBM. (See Appendix G.3.). The training dataset consists of $N_{tr}=40$ trajectories with linearly spaced Reynolds numbers $Re\in\beta_{tr}=[100,200]$, and the test dataset contains $N_{te}=80$ trajectories where 40 trajectories have $Re\in\beta_{tr}$ (parameter interpolation) and another 40 trajectories have $Re\in\beta_{te}=[200,300]$ (parameter extrapolation). In the paper, we report the test errors for parameter interpolation and parameter extrapolation settings separately.
@@ F.1 Implementation
As pointed out in Appendix B.2, we used the JAX implementation of SciPy and NumPy linear algebra utilities for differentiable integration with the training pipeline, support of multi-GPU execution, and easy batch operations. Moreover, as noted in Appendix B.2, the Jacobian matrix is computed using forward-mode AD, which is accessible through JAX’s “jacfwd” function transformation. Finally, the $optimizer$ and $ODESolve$ function calls in Algorithms 1 and 3 are provided by Optax and Diffrax libraries implemented in JAX.
@@ F.5 Training and inference settings
Appendix G Extended results
G.1 Mean squared errors
In the main text, we reported the root normalized errors for our experiments. In cases where the scale of the solutions decays over time (Diffusion, N-S, and KdV), none-normalized error measures such as mean squared error (MSE) may indicate a decreasing error over time, while normalized errors penalize the model for not adapting to the changing scale of the solutions and indicate a growing error. Nevertheless, many other works report results in MSE. For completeness, the mean squared errors for all experiments and both training and test sets are reported here in Tables 7 and 8.
@@ G.1 Mean squared errors
[…]
@@
[…]
@@
[…]
@@
[…]
G.2 Error over time

### v2

@@ Abstract
Reduced-order modeling (ROM) of time-dependent and parameterized differential equations aims to accelerate the simulation of complex high-dimensional systems by learning a compact latent manifold representation that captures the characteristics of the solution fields and their time-dependent dynamics. Although high-fidelity numerical solvers generate the training datasets, they have thus far been excluded from the training process, causing the learned latent dynamics to drift away from the discretized governing physics. This mismatch often limits generalization and forecasting capabilities. In this work, we propose Physics-informed ROM ($\Phi$-ROM) by incorporating differentiable PDE solvers into the training procedure. Specifically, the latent space dynamics and its dependence on PDE parameters are shaped directly by the governing physics encoded in the solver, ensuring a strong correspondence between the full and reduced systems. Our model outperforms state-of-the-art data-driven ROMs and other physics-informed strategies by accurately generalizing to new dynamics arising from unseen parameters, enabling long-term forecasting beyond the training horizon, maintaining continuity in both time and space, and reducing the data cost. Furthermore, $\Phi$-ROM learns to recover and forecast the solution fields even when trained or evaluated with sparse and irregular observations of the fields, providing a flexible framework for field reconstruction and data assimilation. We demonstrate the framework’s robustness across various PDE solvers and highlight its broad applicability by providing an open-source JAX implementation that is readily extensible to other PDE systems and differentiable solvers, available at https://phi-rom.github.io.
@@ 2.3 Reduced-ordered temporal evolution
$\tilde{\alpha}^{t_{0}}=\hat{\alpha}^{t_{0}}=D_{\theta}^{\dagger}(\mathbf{u}^{t_{0}},\mathcal{X})\qquad\tilde{\alpha}^{T_{*}}=\tilde{\alpha}^{t_{0}}+\int_{t_{0}}^{T_{*}}\Psi(\tilde{\alpha}^{t})dt$ (6)
@@ 3.1 Training objective
In Eq. (10), $\lambda$ controls the regularization effect of the dynamics loss. Choice of $\lambda$ depends on the sensitivity of $\mathcal{S}$ to the noise and errors in the reconstructed fields. We chose $\lambda=0.5-0.8$ depending on the solvers (see Appendix F.3 and G for the choice of $\lambda$). See Appendix B for the detailed training procedure.
@@ 4.1 Experimental settings
For each problem, models are trained on $M_{tr}$ trajectories of $T_{tr}$ time steps and evaluated on $M_{te}$ trajectories to assess forecasting and generalization performance. Evaluation is performed by forecasting solutions for unseen initial conditions or parameters in $\mathbf{U}_{te}$ for $T_{tr}$ (interpolation) and $T_{te}$ (extrapolation) time steps. Among the problems, Burgers’ and LBM are parameterized by their source term and Reynolds number, respectively, while the others vary in their initial conditions. Please refer to Appendix F for a detailed description of our experimental settings and Appendix H for extended results.
@@ 4.3 Forecasting performance on full grid
Full training
$\mathcal{X}_{tr}=\mathcal{X}_{\mathcal{S}}=\mathcal{X}_{te}$
@@
Sparse training
$|\mathcal{X}_{tr}|=5\%|\mathcal{X}_{\mathcal{S}}|,\,\mathcal{X}_{te}=\mathcal{X}_{\mathcal{S}}$
@@ 5 Conclusion
We introduced $\Phi$-ROM, a physics-informed ROM that embeds the governing physics in the reduced latent space by learning the dynamics directly from a differentiable PDE solver. We showed that compared to other physics-informed strategies and data-driven methods for constructing ROMs, $\Phi$-ROM generalizes better to unseen dynamics induced by new initial conditions or parameters, forecasts beyond the training temporal horizon, and recovers the full solution fields when trained or tested with partial observations. The model provides accelerated forecasting within a reduced space and benefits from a mesh-free construction that is continuous in space and time. Furthermore, $\Phi$-ROM adheres to the convergent discretization schemes of numerical PDE solvers when learning the temporal dynamics, making it robust to various physical phenomena compared to other physics-informed approaches.
Despite the promising results, we acknowledge that requiring access to differentiable PDE solvers limits $\Phi$-ROM’s immediate applicability, although ongoing developments in differentiable programming offer promising avenues. The inclusion of the PDE solver in the training process also increases the training cost compared to purely data-driven methods. While we believe the superior data efficiency of $\Phi$-ROM justifies the increased training costs (as demonstrated in Fig. 4), tractable physics-informed training for large-scale 3D problems would require further study and optimization of the dynamics loss and its components. Additionally, although $\Phi$-ROM was successfully trained with sparse data, extending the framework to data assimilation settings with limited sensor measurements is a compelling area for further exploration. Finally, this study focused on PDEs with first-order time derivatives, and extending to higher-order time-dependent (e.g. wave equations) or time-independent PDEs is left for future research.
@@ Appendix A Related works
In contrast to the aforementioned works, CROM [21] evolves the solution directly in the physical space using the exact form of the continuous governing PDE (similar to PINNs) at the inference time. Although CROM enforces the underlying equations, making it more interpretable than purely data-driven approaches, its computational savings do not stem from a reduced latent representation of a nonlinear manifold. Instead, it relies on the use of a limited number of sub-sampled grid points that are reconstructed and used to calculate the temporal derivatives of the PDE using automatic differentiation. The time derivatives are then used to find the solution at the next time step which is subsequently encoded to its corresponding latent code. However, due to the errors in the gradients when taking the subsampled time derivatives, CROM is not always able to perform subsampling and instead uses the discretized (e.g. finite-differences) PDE on the full grid to compute the time derivatives. As such, CROM essentially solves the full-order PDE in such scenarios.
Our proposed framework of $\Phi$-ROM adopts a similar setup to DINo [13], but benefits from a physics-informed training strategy. $\Phi$-ROM directly trains the dynamics network with time derivatives of the reconstructed fields, obtained from a differentiable solver, in contrast to DINo, which relies on time-integration during training using neural ODEs to evolve the latent coordinates. We also leverage hyperreduction [22] for fast and efficient training of the dynamics network in $\Phi$-ROM, unlike other similar approaches that also rely on the time derivatives of the reconstructed fields [12, 19]. Moreover, $\Phi$-ROM differs substantially from CROM in the sense that it enforces the governing physics in the latent space and evolves the solution directly in the latent space during the training as opposed to the full physical space.
@@ B.1 Training procedure
[…]
@@ B.2 Dynamics loss and hyper-reduction
2:  $\hat{\mathbf{u}}\leftarrow D_{\theta}(\alpha,\mathcal{X}_{\mathcal{S}})$ $\triangleright$ Reconstruct the field on the solver grid ($\hat{\mathbf{u}}\in\mathbb{R}^{N\times m},N=\left|X_{\mathcal{S}}\right|$)
3:  $\dot{\hat{\mathbf{u}}}\leftarrow(\hat{\mathbf{u}},\mathcal{S})$ $\triangleright$ Compute time derivative, see Appendix B.4
4:  $\mathcal{X}^{\downarrow}_{\mathcal{S}}\sim\mathcal{X}_{\mathcal{S}}$ $\triangleright$ Randomly sample $\gamma N$ coordinates (where $N=|\mathcal{X}_{\mathcal{S}}|$)
5:  $\dot{\hat{\mathbf{u}}}^{\downarrow}\leftarrow\dot{\hat{\mathbf{u}}}|_{\mathcal{X}^{\downarrow}_{\mathcal{S}}}$ $\triangleright$ Sample time derivatives at $\mathcal{X}_{\mathcal{S}}^{\downarrow}$ and flatten ($\dot{\hat{\mathbf{u}}}^{\downarrow}\in\mathbb{R}^{\gamma Nm}$)
6:  $J^{\downarrow}\leftarrow\nabla_{\alpha}D_{\theta}(\alpha,\mathcal{X}^{\downarrow}_{\mathcal{S}})$ $\triangleright$ Jacobian via forward mode AD ($J^{\downarrow}\in\mathbb{R}^{\gamma Nm\times k}$)
7:  $\dot{\alpha}^{*}\leftarrow\operatorname*{\arg\!\min}_{\dot{\alpha}}\left\|\dot{\hat{\mathbf{u}}}^{\downarrow}-J^{\downarrow}\dot{\alpha}\right\|$ $\triangleright$ Solve least-squares ($\dot{\alpha}^{*}\in\mathbb{R}^{k}$)
8:  return $\ell(\Psi_{\phi}(\alpha),\dot{\alpha}^{*})$ $\triangleright$ Return dynamics loss
@@ Appendix C Inference details
[…]
@@
[…]
@@ Appendix D Detailed description of the datasets
Diffusion $100$ $32$ $16$ $25$ $200$ $42\times 42$ 11
Burgers’ 88 99 $100$ $200$ $256$ 11
N-S $256$ $64$ $16$ $50$ $200$ $64\times 64$ 22
KdV $512$ $64$ $16$ $40$ $80$ $64\times 64$ 11
LBM $40$ $40+40$ 44 $100$ $125$ $180\times 82$ 22
@@ D.5 Lattice Boltzmann equation
We used XLB in 2D to simulate flow around a circular cylinder with the D2Q9 lattice, BGK collision operator, and suitable boundary conditions representing the flow inlet and outlet, and the no-slip boundaries for a set of Reynolds numbers, $Re=UD/\nu$ where $U$ is the mean speed of the incoming flow, $D$ is the diameter of the cylinder, and $\nu$ is the kinematic viscosity. For sufficiently large Reynolds numbers, this configuration leads to the formation of the Kármán vortex street caused by vortex shedding. We generate the data on a $180\times 82$ grid with Reynolds numbers ranging from $100$ to $200$ and simulate for $4000$ solver steps. We discard the data from the first $3000$ steps of the solver when the flow is still forming. Furthermore, since LBM makes very small steps, we save the remaining simulated data only every $5$ steps. As such, each trajectory has $200$ time steps, where $T_{tr}=100$ are used for training and the rest are used for testing. However, we report the time extrapolation errors in the paper for only $T_{te}=125$ as both $\Phi$-ROM and DINo fail to capture the long-horizon dynamics of LBM. (See Appendix H.3.). The training dataset consists of $N_{tr}=40$ trajectories with linearly spaced Reynolds numbers $Re\in\beta_{tr}=[100,200]$, and the test dataset contains $N_{te}=80$ trajectories where 40 trajectories have $Re\in\beta_{tr}$ (parameter interpolation) and another 40 trajectories have $Re\in\beta_{te}=[200,300]$ (parameter extrapolation). In the paper, we report the test errors for parameter interpolation and parameter extrapolation settings separately.
@@ F.1 Implementation
As pointed out in Appendix B.2, we used the JAX implementation of SciPy and NumPy linear algebra utilities for differentiable integration with the training pipeline, support of multi-GPU execution, and easy batch operations. Moreover, as noted in Appendix B.2, the Jacobian matrix is computed using forward-mode AD, which is accessible through JAX’s “jacfwd” function transformation. Finally, the optimizeroptimizer and $ODESolve$ function calls in Algorithms 1 and 3 are provided by Optax and Diffrax libraries implemented in JAX.
@@ F.5 Training and inference settings
Appendix G On the regularizing effect of $\lambda$
The hyperparameter $\lambda$ balances the two loss terms in the training objective. Note that while the latent manifold is constructed by $L_{rec}$, it is regularized by $L_{dyn}$ in order to enforce the PDE. Consequently, smaller $\lambda$ increases the regularization effect of $L_{dyn}$ on the decoder and the latent manifold, resulting in a higher reconstruction loss. At the same time, a very large choice of $\lambda$ (>0.9>0.9) may result in a disparity between the actual latent dynamics and what the dynamics network $\Psi$ learns as the latent dynamics. Thus, we suggested the 0.5 to 0.8 range for the choice of $\lambda$. A larger $\lambda$ may improve the forecast error at very early time-steps in exchange for higher error accumulation, while smaller values may deteriorate the reconstructions and result in wrong time derivatives from the solver during the training. In addition, note that as we use normalized errors for both loss functions, the losses would often be in the same order of magnitude when $\lambda$ is close to 0.5 and decrease with similar rates. Below, we further explain our choice of $\lambda$ in the case of Burgers’ and LBM experiments.
G.1 Burgers’ ablation study
To better demonstrate the effect of $\lambda$, we trained Burgers’ with $\lambda=0.1,0.5,0.9$ and report the results in Table 7. As discussed above, increasing $\lambda$ decreases the regularization effect and thus results in smaller reconstruction loss, but fails to minimize the dynamics loss. Conversely, a small $\lambda$ increases the reconstruction error in a way that the dynamics learned from the solver may be wrong or noisy.
Table 7: Loss and error values for Burgers’ with different choices of $\lambda$.
$\lambda$ $L_{rec}$ $L_{dyn}$ Error $T_{tr}$ Error $T_{te}$
0.9 0.002 0.081 0.137 0.150
0.5 0.019 0.013 0.021 0.028
0.1 0.176 0.011 0.459 0.613
G.2 LBM case
In our experiments, $\lambda=0.5$ worked well for all problems, except for LBM, which has a more complex geometry because of the discontinuity at the cylinder. Due to the imperfect reconstruction of the cylinder, we observed noise and unexpected artifacts in the time derivatives given by the PDE solver. To alleviate this issue, we increased $\lambda$ to achieve a smaller reconstruction loss, which helped with stable training of the model and improving its generalization.
Appendix H Extended results
H.1 Mean squared errors
In the main text, we reported the root normalized errors for our experiments. In cases where the scale of the solutions decays over time (Diffusion, N-S, and KdV), non-normalized error measures such as mean squared error (MSE) may indicate a decreasing error over time, while normalized errors penalize the model for not adapting to the changing scale of the solutions and indicate a growing error. Nevertheless, many other works report results in MSE. For completeness, the mean squared errors for all experiments and both training and test sets are reported here in Tables 8 and 9.
@@ H.1 Mean squared errors
$\Phi$-ROM $6.0\text{e-}4$ $\mathbf{8.2\textbf{e-}4}$ $\mathbf{2.9\textbf{e-}5}$ $\mathbf{3.8\textbf{e-}5}$ $2.5\text{e-}3$ $2.6\text{e-}3$ $\mathbf{9.6\textbf{e-}3}$ $\mathbf{1.1\textbf{e-}2}$
DINo $\mathbf{3.6\textbf{e-}4}$ $8.8\text{e-}4$ $9.4\text{e-}5$ $1.4\text{e-}4$ $2.6\text{e-}3$ $2.7\text{e-}3$ $6.4\text{e-}2$ $1.0\text{e-}1$
PINN-ROM $5.3\text{e-}4$ $8.2\text{e-}4$ $6.0\text{e-}5$ $8.0\text{e-}5$ $7.3\text{e-}2$ $7.5\text{e-}2$ $1.4\text{e+}0$ $1.3\text{e+}0$
FD-CROM $1.0\text{e-}3$ $1.7\text{e-}3$ $5.1\text{e-}3$ $6.3\text{e-}3$ $\mathbf{9.8\textbf{e-}6}$ $\mathbf{1.2\textbf{e-}5}$ $5.7\text{e-}2$ $8.1\text{e-}2$
AD-CROM $4.0\text{e-}4$ $9.3\text{e-}4$ $4.5\text{e-}4$ $6.0\text{e-}4$ $1\text{e-}1 1.1$$1\text{e-}2$ $4.2\text{e-}1$ $4.3\text{e-}2$
@@
$\Phi$-ROM $2.4\text{e-}4$ $1.7\text{e-}3$ $2.2\text{e-}4$ $2.0\text{e-}3$ $6.3\text{e-}3$ $7.0\text{e-}3$ $1.6\text{e-}2$ $1.7\text{e-}2$ $8.5\text{e-}6$ 5.0e-5 $5.9\text{e-}5$ $1.4\text{e-}4$
DINo $1.1\text{e-}4$ $1.9\text{e-}2$ $5.1\text{e-}3$ $3.0\text{e-}2$ $1.1\text{e-}3$ $2.6\text{e-}2$ $1.1\text{e-}2$ $3.6\text{e-}2$ $3.6\text{e-}7$ 5.3e-5 $2.6\text{e-}3$ $3.4\text{e-3}$
@@
$\Phi$-ROM $2.5\text{e-}4$ $1.7\text{e-}3$ $2.3\text{e-}4$ $2.0\text{e-}3$ $6.3\text{e-}3$ $7.0\text{e-}3$ $1.7\text{e-}2$ $1.7\text{e-}2$ $8.4\text{e-}6$ $5.0\text{e-}5$ $4.8\text{e-}4$ $1.8\text{e-}3$
DINo $1.4\text{e-}4$ $1.9\text{e-}2$ $5.1\text{e-}3$ $3.3\text{e-}2$ $1.2\text{e-}3$ $2.6\text{e-}2$ $1.1\text{e-}2$ $3.6\text{e-}2$ $3.7\text{e-}7$ $5.3\text{e-5}$ $2.4\text{e-}3$ $3.2\text{e-}3$
@@
$\Phi$-ROM $3.5\text{e-}4$ $1.8\text{e-}3$ $9\text{e-}4 2.2$$2\text{e-}3$ $3\text{e-}2 5.5$$\text{e-}2 5.5$$5\text{e-}2$ $5.7\text{e-}2$ $8.5\text{e-}6$ $5.0\text{e-}5$ $4.8\text{e-}4$ $1.8\text{e-}3$
DINo $6.3\text{e-}4$ $2.0\text{e-}2$ $5.1\text{e-}3$ $3.2\text{e-}2$ $3.2\text{e-}2$ $4.6\text{e-}2$ $3.1\text{e-}2$ $4.2\text{e-}2$ $3.7\text{e-}7$ $5.3\text{e-}5$ $2.1\text{e-}3$ $3.0\text{e-}3$
H.2 Error over time